OpenStudy (anonymous):

HELP ASAP 1/10 - (4/5-3/4) - (2/15 multiplied by 3/8)

OpenStudy (anonymous):

the answer is zero.

OpenStudy (anonymous):

can you please show the work

OpenStudy (anonymous):

sure. 1/10-((16-15)/20)-1/20 =1/10-1/20-1/20 =1/10-1/10 =0 Am I becoming foolish and missing something? I think this question is a child's play.